Show that a triangular matrix is normal if it is diagonal


1
Expert's answer
2021-02-09T01:35:04-0500

Answer:


Given:

A triangular matrix is normal if it is diagonal.


For n=2, Let A= "\begin{bmatrix}
 a & 0 \\
 b & c
\end{bmatrix}" such that

AA=AA= "\\begin{bmatrix}\n \\overline a & b \\\\\n b & \\overline c\n\\end{bmatrix}\n\n\u200b""\begin{bmatrix}
 a & 0 \\
 b & c
\end{bmatrix}" -"\begin{bmatrix}
 a & 0 \\
 b & c
\end{bmatrix}""\\begin{bmatrix}\n \\overline a & \\overline b \\\\\n 0 & \\overline c\n\\end{bmatrix}\n\n\u200b"

= "\begin{bmatrix}
 ||a||^2+||b||^2−||a||^2 & \overline bc−a \overline b \\
 \overline cb−b \overline a & |c||^2−(||b||^2+||c||^2)
\end{bmatrix}"


= "\begin{bmatrix}
 b||^2 & \overline bc−a \overline b \\
 \overline cb−b \overline a & −||b||^2
\end{bmatrix}"


So b=0 and A is diagonal ,if the result is true.For n≥2, Let A= "\begin{bmatrix}
 T & 0 \\
 v & a
\end{bmatrix}" where T is a n x n triangular matrix,v a 1 x n matrix and a is a complex number . Since AA=AA we have,


"\begin{bmatrix}
 0 & 0 \\
 0 & 0
\end{bmatrix}"= "\begin{bmatrix}
 T^* & v* \\
 0 & \overline a
\end{bmatrix}""\begin{bmatrix}
 T & 0 \\
 v & a
\end{bmatrix}"-"\begin{bmatrix}
 T & 0 \\
 v & a
\end{bmatrix}" "\begin{bmatrix}
 T^* & v* \\
 0 & \overline a
\end{bmatrix}"


=  "\begin{bmatrix}
 T^*T + vv*& v*a \\
 \overline a v & |a|^2
\end{bmatrix}"-  "\begin{bmatrix}
 TT^* & Tv* \\
 vT* & |v|^2 +|a|^2
\end{bmatrix}"


=  "\begin{bmatrix}
 T^*T- TT*+vv* & v*a-Tv* \\
 \overline a v-vT & -|v|^2
\end{bmatrix}"


Hence v=0 and T is normal .since T is lower triangular,then T is normal if A is diagonal.